### v1

# DiTSE: High-Fidelity Generative Speech Enhancement via Latent Diffusion Transformers

Abstract
Real-world speech recordings suffer from degradations such as background noise and reverberation. Speech enhancement aims to mitigate these issues by generating clean high-fidelity signals. While recent generative approaches for speech enhancement have shown promising results, they still face two major challenges: (1) content hallucination, where plausible phonemes generated differ from the original utterance; and (2) inconsistency, failing to preserve speaker’s identity and paralinguistic features from the input speech. In this work, we introduce DiTSE (Diffusion Transformer for Speech Enhancement), which addresses quality issues of degraded speech in full bandwidth. Our approach employs a latent diffusion transformer model together with robust conditioning features, effectively addressing these challenges while remaining computationally efficient. Experimental results from both subjective and objective evaluations demonstrate that DiTSE achieves state-of-the-art audio quality that, for the first time, matches real studio-quality audio from the DAPS dataset. Furthermore, DiTSE significantly improves the preservation of speaker identity and content fidelity, reducing hallucinations across datasets compared to state-of-the-art enhancers. Audio samples are available at: http://hguimaraes.me/DiTSE
Index Terms:
Diffusion Models, Diffusion Transformers, Speech Enhancement, Dereverberation, Generative Models.I Introduction
Real-world recorded speech is often degraded by acoustic conditions such as noise, reverberation, and equalization issues. Speech enhancement methods aim to improve such signals’ perceptual quality and intelligibility [1], with useful applications to audio content creation and telecommunications. Recent advances in deep learning have established new performance standards for speech enhancement. While existing methods can alleviate some quality issues, achieving clean high-fidelity audio remains a challenge [2].
The conventional enhancement methods based on deep neural networks learn a deterministic many-to-one mapping via a regressive objective, which maps input signals that have been variously degraded from a clean signal to the corresponding ground-truth clean signal. However, the problem becomes multi-modal in challenging scenarios where degradations, such as reverberation, create ambiguity in determining the target phase of the clean signal. Therefore, the conventional methods often experience mode collapse as they predict an averaged version of the many plausible signals, resulting in artifacts and distortion in their enhanced results.
Generative methods directly model the distribution of clean speech, avoiding mode collapse and enabling effective utilization of speech priors. They aim to generate a plausible version of many possible high-fidelity clean signals that preserve spoken content and speaker traits of the input degraded audio. Recently, discrete generation methods for speech enhancement, often using neural codec tokens, have shown significant improvement of audio clarity and cleanness over conventional methods. They employ language modeling [3, 4] or masked language modeling [5, 6] on tokens while conditioning on the input signal. In particular, Genhancer [5] incorporates speech features to improve speaker consistency and content accuracy. However, the token spaces limited by bit-rates often miss nuances in spoken content, thus compromising the naturalness of enhanced voices. Moreover, the costly sampling processes at inference time make the methods impractical for many real-world applications.
On the other hand, diffusion-based methods operate in a continuous space (i.e., sample or latent representation), stemming from the common concept of non-equilibrium thermodynamics of learning to reverse noise back to the data distribution iteratively. Particularly, SGMSE+ [7] and CDiffuSE [8] interpolate noisy speech with Gaussian noise via the forward diffusion process, aiming to reduce both the Gaussian noise and environmental degradations during the reverse process. Alternatively, UNIVERSE [9] and UNIVERSE++ [10] model speech generation conditioning on the high-level features extracted from the input audio. Moreover, StoRM [11] leverages a diffusion process to refine the outputs from other enhancement systems via regeneration, combining benefits from both conventional and generative approaches.
We propose DiTSE (Diffusion Transformer for Speech Enhancement), a high-fidelity full-bandwidth enhancer using a latent diffusion transformer model. DiTSE performs diffusion in a learnt audio latent space with 32 sampling steps, effectively reducing computational burden over previous diffusion-based approaches. Comprehensive objective and subjective evaluations show that DiTSE better maintains speaker traits and reduces content hallucination compared to prior art, while achieving state-of-the-art audio quality that, for the first time, matches real studio-quality audio from the DAPS dataset [12]. We additionally introduce the technique of post-hoc quantization to further improve audio fidelity.
II Proposed Method
We present the three main components of DiTSE: (1) a variational autoencoder; (2) a diffusion transformer model; (3) a conditioner network that guides the generation process with pre-trained and learned speech features. In addition, we introduce a post-hoc process to refine the audio quality using discrete generation.
II-A Preliminaries on Diffusion Models
Diffusion-based models have drawn attention due to their generation capabilities in domains such as images [13] and audio [14, 15]. Diffusion models employ a forward process that adds Gaussian noise to clean input data [16], which can be expressed in marginal , where are hyperparameters from the noise scheduler [17] at the timestep . The objective is to learn to reverse this process via a neural network. However, instead of directly trying to predict the data reconstruction , it has been empirically found that other objectives are more stable to produce high-quality outputs. In this work, we use the v-prediction reparametrization [18], where the model predicts the velocity vector, , where . For the interested reader in the formulation, we refer to [19, 18, 17].
II-B Variational Autoencoder
We learn a variational autoencoder on publicly available audio data (i.e., speech, music and sound events) to encode audio signals of sample rate into latent representations of 64 dimensions at frame rate. The VAE uses the network structure of a neural audio codec - DAC [20], with the quantization layers replaced by a variational bottleneck. To ensure high-quality signal reconstruction, the VAE model also adopts the GAN setup of DAC, adversarially trained with multi-period waveform discriminators and multi-scale multi-band complex spectrogram discriminators. The learned latent space, which is compressed in representation and regularized in distribution, facilitates computationally efficient generation and high-quality decoding into audio signals.
II-C Latent Diffusion Transformer (DiT) Network
While U-Net architectures commonly serve as backbones for diffusion-based speech enhancement methods [21, 7], recent advancements in audio generation [22, 23] have shown that Transformer-based backbones offer significant advantages in scalability and robustness. Herein, we adopt Diffusion Transformer (DiT) [24] on the latent representations from the VAE, leveraging established practices and strategies from the vision domain for training and inference. The model conditions on the timestep embedding of the current diffusion step via adaptive layer normalization following common practices. Other conditioning information (discussed below) is provided to the network either via cross-attention or via concatenation with the diffusion latent along the feature dimension. We introduce 50% probability of prefix prompting in training, where a random number of frames at the beginning of the diffusion latent is unmasked with the ground-truth clean latent. During enhancement of long audio, we provide a previously enhanced audio segment as a prompt to select a consistent generation mode across windows of windowed generation.
II-D Robust Conditioning Features
The conditional speech generation should recover clean audio that faithfully preserves the original spoken content and paralinguistic features of the input audio. We take inspiration from the prior work Genhancer [5] where the generative model conditions on two kinds of speech features: self-supervised learned (SSL) features extracted from the input audio, and a learned representation of the input audio.
Robust SSL features. Leveraging large-scale pre-training on speech data, SSL features provide strong semantic guidance for spoken content, thus reducing content hallucinations in enhancement. This work experiments with WavLM [25] and RobustDistiller [26, 27]. The DiT network cross-attends to the learned weighted sum of feature maps of the SSL model, as the features differ in temporal resolutions from the latent.
Conditioner Network. As indicated in Figure 1, a feed-forward network using Conformer [28] architecture takes in the VAE latent and transforms it to a cleaner representation that approximates the clean latent. We jointly train the Conditioner network with the main diffusion network, using L1-loss against the ground-truth clean latent. Our preliminary study shows that stacking this pre-enhanced latent with the diffusion latent can exploit the explicit temporal alignment between the two and thus works more effectively than conditioning via cross-attention. We can view the design as if the diffusion network observes two perspectives of the clean latent at each diffusion step – one corrupted by Gaussian noise, and one corrupted by acoustic conditions translated to the latent space.
Auxiliary Timestep Embedding. Building on the timestep conditioning described earlier, we additionally concatenate the timestep embedding with the input of the diffusion network. The practice enhances the model’s capability to adaptively fuse the diffusion latent and the pre-enhanced latent at different diffusion steps. By intuition, the pre-enhanced latent provides a direct path for structuring a speech envelope in the early diffusion steps, but introduces distractions caused by its own errors and artifacts in the later steps. Without this auxiliary timestep embedding, we noticed from qualitative assessments that residuals of the noise and reverb in the input audio are more likely to be carried over to the final output.
II-E Post-hoc Quantization
The design above achieves good content accuracy and speaker preservation across audio samples, but displays instability in acoustic quality compared to discrete generation methods (e.g., Genhancer [5]). We hypothesize that: (1) the distance in the latent space does not correlate well with the perceptual difference, and thus, small errors could lead to perceptually significant acoustic quality degradations; (2) the diffusion trajectory is sensitive to the starting noise and sampling configurations, and is prone to accumulating errors for out-of-domain data. While prior work [10] addressed perceptual issues by fine-tuning the final diffusion steps, we propose to leverage discrete generation methods post-hoc for their benefits in audio fidelity. Specifically, we decode the predicted clean VAE latent into audio and use the discrete neural codec DAC [20] to reconstruct the audio, eliminating audible artifacts and regularizing harmonics of the speech from the latent diffusion generation.
III Experiments
III-A Experiment Setup
Training Datasets. Our models are trained on publicly available datasets that simulate real-world acoustic conditions. Note that none of these datasets are used for evaluation. For clean speech, we use LibriTTS-R [29], a quality-improved 585-hour dataset, further upsampled to sample rate via bandwidth extension [30]. Noise samples are uniformly sampled from three datasets: (1) SFS-Static-Dataset [31], (2) TAU Urban Audio-Visual Scenes 2021 dataset [32], and (3) the DNS Challenge [33]. Reverberation is simulated using room impulse responses (RIRs) from the OpenSLR28 dataset [34], MIT IR Survey [35], and EchoThief [36]. To generate the pairs of synthetic training data, we employ a data simulation and augmentation procedure based on previous work [5, 37]. A speech signal is first convolved with a RIR and then mixed with one or two noise samples at a signal-to-noise ratio (SNR) uniformly sampled between -10 dB and 20 dB. The signal also undergoes random equalization and bandwidth limitation.
Hyperparameters and Hardware. The DiT architecture consists of 12 Transformer layers with 8 attention heads. We evaluate two configurations: Base (335M parameters, 1024 channels, conditioned on WavLM features) and Tiny (112M parameters, 512 channels, conditioned on RobustDistiller features [26, 27], distilled from WavLM). Sinusoidal positional encodings are used across all attention layers. The conditioner branch uses a Conformer model of 12 layers, each with an input size, feed-forward dimension, and kernel size of 256, 1024, and 31, respectively. We apply Classifier-Free Guidance [38] during training with a probability of 10% to encourage the model to effectively leverage each conditioning feature. We train our models on eight NVIDIA A100 GPUs for 400k steps using a batch size of 128 5 seconds together with the AdamW optimizer [39]. The learning rate is linearly warmed up from 0 to over the first 10% training steps before decaying to in subsequent steps. We chose the cosine noise scheduler over others after qualitative comparison of the generated results. We adopt steps at inference as additional steps yield minimal improvements, and use dpmpp-3m-sde sampler from k-diffusion111https://github.com/crowsonkb/k-diffusion library.
| Experiment | WER | WB-PESQ | ESTOI | DNSMOS |
| Input | 5.03 | 1.43 | 66.83 | 2.49 |
| Ablation Studies | ||||
| Baseline | 4.29 0.71 | 2.31 0.01 | 84.24 0.06 | 3.32 0.01 |
[…]
| (+) Post-hoc Quantization | 4.01 0.61 | 2.35 0.01 | 85.15 0.09 | 3.32 0.01 |
| Content [WER ] | Speaker [MOS ] | Quality [MOS ] | |||||||
| DAPS (16 kHz) | VBD (16 kHz) | EARS (16 kHz) | VBD (48 kHz) | EARS (48 kHz) | DAPS (48 kHz) | AQECC (16 kHz) | DEMO (16 kHz) | VBD (48 kHz) | |
| Input | 5.03 | 5.70 | 15.08 | 4.84 0.02 | 4.89 0.02 | 1.74 0.04 | 2.74 0.06 | 1.93 0.05 | 1.84 0.05 |
| Clean | — | — | — | 4.41 0.04 | 4.18 0.05 | 4.30 0.04 | — | — | 3.87 0.05 |
| VAE Reconst. | 0.87 | 1.23 | 3.34 | 4.28 0.04 | 4.10 0.05 | 4.25 0.04 | — | — | 3.54 0.05 |
| HiFi-GAN-2 [37] | 9.16 | 6.63 | 23.67 | 4.09 0.05 | 3.16 0.07 | 3.63 0.05 | 3.70 0.05 | 3.33 0.06 | 3.77 0.05 |
| Miipher [40] | 7.64 | 12.02 | 32.32 | 3.24 0.06 | 2.68 0.06 | 3.40 0.05 | 3.53 0.05 | 3.15 0.05 | 3.42 0.06 |
| Genhancer [5] | 6.04 | 6.47 | 26.33 | 4.03 0.05 | 3.27 0.07 | 4.08 0.05 | 3.89 0.04 | 3.78 0.05 | 4.03 0.05 |
| SGMSE+ [7] | 6.69 | 10.32 | 33.60 | 3.87 0.06 | 3.03 0.06 | 3.51 0.05 | 3.71 0.04 | 3.33 0.05 | 3.59 0.06 |
| StoRM [11] | 11.27 | 9.86 | 45.87 | 3.54 0.06 | 2.75 0.06 | 2.62 0.05 | 3.61 0.05 | 2.53 0.06 | 2.94 0.05 |
| (Ours) DiTSE-Tiny | 8.58 | 9.55 | 38.21 | 3.89 0.06 | 2.92 0.07 | 4.01 0.04 | 3.83 0.05 | 3.30 0.05 | 3.63 0.05 |
| (Ours) DiTSE-Base | 3.56 | 4.93 | 26.42 | 4.20 0.04 | 3.45 0.06 | 4.34 0.04 | 3.88 0.05 | 3.61 0.05 | 3.91 0.05 |
| (Ours) DiTSE-Base + Post | 3.71 | 5.39 | 26.72 | 4.27 0.04 | 3.50 0.06 | 4.32 0.04 | 3.97 0.04 | 3.77 0.05 | 4.00 0.05 |
III-B Ablation Study
We investigate the design choices of the proposed DiTSE model on the DAPS dataset [12], which consists of pairs of studio-quality speech recordings and their degraded versions under real-world acoustic conditions. Our analysis utilizes a subset of DAPS as used by prior work [5] – one female voice (f10) and one male voice (m10) reading a 2-minute script (script5). Table I presents the objective scores of Word Error Rate (WER)222Calculated between the enhanced audio and the clean reference audio transcribed by Whisper-Large-v3-v20240930 (https://github.com/openai/whisper), WideBand Perceptual Evaluation of Speech Quality (WB-PESQ), Extended Short-Time Objective Intelligibility (ESTOI), and the non-intrusive metric DNSMOS [41]. Given the stochastic nature of generative approaches, we perform inference five times for each model using different random seeds and report the statistics across all the runs.
Our Baseline approach, a DiT model without concatenating timestep embeddings or a Conditioner network, achieved notable content preservation. Incorporating additional timestep embeddings at the input improves acoustic quality. Introducing a Conditioner network further improves WER, with the masking formulation (i.e., predicting a non-negative mask) showing slightly better stability than the mapping formulation (i.e., directly predicting the features) in the latent space. Meanwhile, the smaller model, with 4.5 times fewer parameters, decreases in content accuracy but retains reasonable overall quality. Lastly, applying the post-hoc quantization to the enhanced result from our full-sized model improves its perceived audio quality, but compromises slightly content fidelity.
III-C Three-axis Evaluations
We compare the proposed methods to state-of-the-art baselines using objective and subjective evaluations. We consider three evaluations axes: (1) content accuracy measured by WER; (2) acoustic quality measured by quality mean-opinion-score (MOS); (3) speaker preservation measured by similarity MOS. We examine the methods on five evaluation sets unseen during training: DAPS Real, AQECC and DEMO, as utilized in prior work [5], each consisting of 100 samples representing diverse real-world scenarios; VoiceBank-DEMAND (VBD) [42] for denoising benchmark; and EARS [43] for challenging denoising and de-reverberation scenarios. For VBD and EARS, we sample twenty distinct speakers in each category with two samples per speaker, ensuring a variety of voices with balanced genders. We compare with the following methods: HiFi-GAN-2 [37], a GAN-based method for high-quality enhancement at ; transcription-free Miipher [40], a generative method which re-synthesizes clean speech from enhanced SSL features at ; Genhancer [5], a discrete token generation method at that is closest to our method in conditioning designs; SGMSE+ [7] at and StoRM [11] at as diffusion baselines. For generative methods, we evaluate the samples of the inference run with the median WER from five random runs. We use audio samples at the maximum sample rates offered by the models and the evaluations sets. We additionally include and down-sampled versions of our results for fair comparisons to the baselines at their sample rates.
We conducted MOS listening tests via Prolific333https://www.prolific.co/. to evaluate audio quality and speaker similarity. For quality, listeners rated audio samples from 1 (Bad) to 5 (Excellent) in consideration of cleanness and naturalness, using a fixed clean DAPS sample as the reference of studio audio quality. For speaker similarity, listeners rated the preservation of the speaker’s voice in an audio sample with respect to the corresponding input audio, using the same scale. Each method received at least 400 ratings for each evaluation set in each test. More details on the listening test setup can be found on our demo page.
Table II reports the results for the three evaluation axes. DiTSE stands out as the only model capable of reducing content error compared to the noisy input on DAPS and VBD, significantly outperforming all the baselines. EARS is particularly challenging as its spoken content is barely intelligible due to severe corruption by noise and reverberation. DiTSE is among the best of the generative methods on EARS, while HiFi-GAN-2 achieves slightly higher content accuracy but lower enhanced quality by manipulating the original waveform. Striking the balance between enabling content restoration and refraining models from hallucinations still remains an open problem under extreme SNR scenarios.
Our methods significantly improve acoustic quality over all the baselines. To the best of our knowledge, DiTSE-Base is the first method that achieves a Quality MOS of 4.34 indistinguishable from the real studio-quality clean audio on DAPS. It also achieves an MOS comparable with Genhancer on AQECC, but under-performs on DEMO and VBD. We pinpoint that the VAE could limit audio quality as the reconstruction of clean audio (VAE Reconst.) displays notable reduction in Quality MOS from the ground-truth clean audio on VBD. Incorporating the post-hoc quantization (DiTSE-Base + Post) improves audio fidelity, achieving state-of-the-art audio quality across all the evaluation sets, while the content remains largely unchanged aside from a few subtle pronunciation variations at difficult spots (<0.5% WER increase).
Lastly, Speaker MOS shows significant advantages of DiTSE in preserving speaker and paralinguistic features of the original speech over baselines. The robust conditioning enables the model to effectively attend to speaker traits in complex acoustic scenarios. Furthermore, generation in the continuous space facilitates the preservation of spoken nuances such as breath and coarticulation, resulting in more natural sounding compared to discrete token generation, while quantization after generation could further enhance fidelity of the voice.
IV Conclusions
We introduce DiTSE, a speech enhancement approach built upon a latent diffusion transformer model. DiTSE effectively preserves spoken content and speaker characteristics by incorporating robust conditioning features. We additionally leverage post-hoc quantization to mitigate audio fidelity issues of diffusion-based approaches. Comprehensive evaluations demonstrate that our proposed approach efficiently achieves state-of-the-art audio quality comparable to real studio quality in 32 generation steps, while significantly improving content accuracy and speaker consistency over previous methods. Diffusion distillation techniques [44] could further reduce the computational costs of our approaches in the future.